OpenStudy (anonymous):

Which of the following constants can be added to x^2 - x to form a perfect square trinomial? 1/4 1/2 1

OpenStudy (anonymous):

example (x + 1)(x + 1) = x^2 + 2x + 1 <----- perfect square trinomial

OpenStudy (anonymous):

x^2 - x divide -1 by 2 gives -1/2 now square -1/2 (-1/2) * (-1/2) = ?

OpenStudy (anonymous):

1/4? @Peach-es1234

OpenStudy (anonymous):

I think so :)

OpenStudy (anonymous):

thank you